Multiply every term in the equation 2x2 + 10x + 12 = 0 by a number other than 0.
Answer by stanbon(75887)

Multiply every term in the equation 2x2 + 10x + 12 = 0 by a number other than 0.
------
Multiply thru by 1/2 to get:
x^2 + 5x + 6 = 0
------
Factor:
(x+3)(x+2) = 0
-----
x = -3 or x = -2
====================
Cheers,
Stan H.
==================